OpenStudy (anonymous):

Which of the following equations is an example of an inverse variation between the varibales x & y? My answer choices are y=x/9, y=9/x, y=x+9, y=9x

OpenStudy (amistre64):

y = k/x is an inverse variation

OpenStudy (anonymous):

@amistre64 would that equate to y=x/9?

OpenStudy (amistre64):

no, x is underneath, not the constant

OpenStudy (anonymous):

@amistre64 ah, ok I misunderstood. so it would be y=9/x

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

@amistre64 thank you for your help!

OpenStudy (amistre64):

youre welcome